OpenStudy (anonymous):

Algebra 2 help? it looks familiar too

OpenStudy (anonymous):

\[\sqrt{80}-3\sqrt{63}-\sqrt{125}-\sqrt{175}\]

OpenStudy (amistre64):

lol, in this case you want to see if you can simplify the radicals

OpenStudy (anonymous):

help with what, simplifying that and into what?

OpenStudy (anonymous):

yeah i simplified \[\sqrt{80}\] to 9

OpenStudy (anonymous):

the rest didnt simplify

OpenStudy (amistre64):

reducing the number to prime factors is a good exercise \[\sqrt{80}-3\sqrt{63}-\sqrt{125}-\sqrt{175}\] \[\sqrt{2*40}-3\sqrt{9*7}-\sqrt{5*25}-\sqrt{25*7}\] \[\sqrt{2*4*10}-3\sqrt{3*3*7}-\sqrt{5*5*5}-\sqrt{5*5*7}\] \[\sqrt{2*2*2*2*5}-3\sqrt{3*3*7}-\sqrt{5*5*5}-\sqrt{5*5*7}\]

OpenStudy (amistre64):

everything that is a doubled is a perfect square \[\sqrt{2*2}\sqrt{2*2}\sqrt{5}-3\sqrt{3*3}\sqrt{7}-\sqrt{5*5}\sqrt{5}-\sqrt{5*5}\sqrt{7}\] \[2*2\sqrt{5}-3*3\sqrt{7}-5\sqrt{5}-5\sqrt{7}\] \[4\sqrt{5}-9\sqrt{7}-5\sqrt{5}-5\sqrt{7}\] now we are back to 2 different radicals that cant be simplified any more

OpenStudy (anonymous):

now is this to the same format of the last question?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ok thanks...fyi all the question in the assignment are different so im going to be asking more questions

OpenStudy (amistre64):

thats fine, the database has lots of room :)